OpenStudy (anonymous):

what is the wavelength of a wave with a frequency of 4.00x10^8

hartnn (hartnn):

frequency = speed of light / wavelength so 4*10^8 = 3* 10^8 / wavelength so wavelength = ?

OpenStudy (anonymous):

umm 1.2x10^17

OpenStudy (anonymous):

is that wrong

hartnn (hartnn):

incorrect wavelength = 3*10^8 / 4*10^8 = ?

OpenStudy (anonymous):

.75

hartnn (hartnn):

yup. thats correct

hartnn (hartnn):

so your wavelength is 0.75 meters